OpenStudy (anonymous):

please help me solve these using u substitution ( please show the workings) 1. e^x(root(1+e^x)) 2.xtan(x^2) 3 sin(rootx)/rootx

OpenStudy (anonymous):

Are you looking to take a derivative?

OpenStudy (anonymous):

My guess would be the integral

OpenStudy (anonymous):

please integrate them through u subs

OpenStudy (anonymous):

OK, which one are u doing tbates?

OpenStudy (anonymous):

the second one I have

OpenStudy (lalaly):

1. u = 1+e^x du = e^x dx so \[\int\limits{\sqrt{u}du}\]

OpenStudy (lalaly):

2. u= x^2 du= 2xdx \[\frac{1}{2} \int\limits{tanu du}\]

OpenStudy (anonymous):

On the second one set u=x^2 du = 2x dx du/2=2xdx Sub in: tan(u)/2 du The integral of tan(u)/2 -(1/2)ln|cos u| + C Plug back in for x -(1/2)ln|cos x^2| + C

OpenStudy (anonymous):

thanx guys and number 3?

OpenStudy (lalaly):

u= sqrtx du = 1/(2sqrtx) \[2 \int\limits{sinu du}\]